17. A net force of 24 N is acting on a 4.0-kg object. Find the acceleration in m/s2
Answer #1

ANSWER.

   17. Calculation of acceleration in meter/second square.

              According to Newton's second law ( F = m * a )

                   Formula for calculate acceleration from a force is

                                F = m * a

                              a = F / m

                 Here

                         Force (N) = 24 N.

                        mass ( Kg) = 4 kg.

       So as per formula

               Acceleration = F / m

                                        = 24 / 4

                                        = 6 meter / second square.